### v2

# Cost-Efficient Continual Learning with Sufficient Exemplar Memory

Abstract
Continual learning (CL) research typically assumes highly constrained exemplar memory resources. However, in many real-world scenarios—especially in the era of large foundation models—memory is abundant, while GPU computational costs are the primary bottleneck. In this work, we investigate CL in a novel setting where exemplar memory is ample (i.e., sufficient exemplar memory). Unlike prior methods designed for strict exemplar memory constraints, we propose a simple yet effective approach that directly operates in the model’s weight space through a combination of weight resetting and averaging techniques. Our method achieves state-of-the-art performance while reducing the computational cost to a quarter or third of existing methods. These findings challenge conventional CL assumptions and provide a practical baseline for computationally efficient CL applications.
1 Introduction
Continual learning (CL) has attracted significant attention as a paradigm enabling machine learning models to adapt to sequential tasks while overcoming catastrophic forgetting of previously acquired knowledge (Wang et al., 2024). A fundamental challenge in CL is the stability-plasticity dilemma (Carpenter & Grossberg, 1987; Mermillod et al., 2013), where a model that maintains stability on previous tasks often struggles to integrate new knowledge, and vice versa. Among various approaches, exemplar-based methods that store representative samples from past tasks have been widely adopted, particularly in class-incremental learning (class-IL) (Masana et al., 2022). Traditionally, these methods have operated under strict exemplar memory constraints, assuming that minimizing memory usage is critical for cost efficiency. However, recent studies suggest that memory limitations may not be as critical as once thought, with computational costs—such as GPU usage—emerging as a more dominant factor (Prabhu et al., 2023; Chavan et al., 2023). For instance, an AWS GPU server equipped with 8 A100 GPUs incurs an hourly cost of approximately \$32.77, whereas storing 1TB of data on Amazon S3 costs roughly \$23.55 per month.
Building on this, we revisit the class-IL setting under more realistic conditions, where exemplar memory constraints are relaxed (i.e., sufficient exemplar memory is available). Our experiments reveal that even the simplest exemplar-based approach—such as naive finetuning (i.e., Replay)—performs comparably to state-of-the-art algorithms while incurring significantly lower training costs (e.g., shorter training time). Motivated by this, we propose a simple yet effective approach that optimizes computational cost while maintaining strong performance. Our method is designed with two key objectives: (1) outperforming other algorithms including Replay, and also (2) achieving computational efficiency, in the setting with sufficient exemplar memory. The core idea is to lower computational overhead by directly operating in the model’s weight space. Specifically, we integrate weight resetting and averaging techniques to accelerate convergence while enhancing both stability and plasticity during CL. Experimental results show that our method not only surpasses competing approaches but also achieves superior performance while requiring only a quarter to a third of the computational cost of state-of-the-art algorithms. Our contributions are summarized as follows:
  • We investigate the impact of relaxed exemplar memory constraints on class-IL algorithms and demonstrate that even naive finetuning performs competitively under these conditions while significantly reducing the computational cost.
  • We propose a novel weight-space manipulation strategy that achieves superior performance while drastically reducing computational overhead compared to state-of-the-art algorithms.
2 Related Works
Continual learning   Continual learning (CL) has been actively studied in various scenarios and methodological categories. Among the three scenarios of CL (Van de Ven & Tolias, 2019), class-incremental learning (class-IL) has been considered the most challenging and has been the most actively studied scenario (Masana et al., 2022). CL algorithms (including class-IL) can be categorized into regularization-based approaches, which penalize changes to important parameters for past tasks (Kirkpatrick et al., 2017; Aljundi et al., 2018; Cha et al., 2020), rehearsal-based approaches, which store and replay exemplars from past tasks (Rebuffi et al., 2001; Cha et al., 2023), and expansion-based approaches, which expand the model’s capacity to balance the trade-off between stability and plasticity (Yan et al., 2021; Wang et al., 2022). Additional approaches focus on addressing classifier bias toward recent tasks while using the exemplars (Wu et al., 2019; Zhao et al., 2020). While exemplar-based methods have demonstrated state-of-the-art performance, they typically rely on strict memory constraints, often limiting memory size to a small percentage of the dataset. Recent studies challenge the necessity of these strict memory constraints, highlighting that the computational cost of maintaining and processing memory—especially GPU usage—can far outweigh storage costs (Prabhu et al., 2023; Chavan et al., 2023; Harun et al., 2023). This shift in perspective opens the door to relaxing memory limits in order to reduce training costs, which is the focus of our work.
Weight space operations Weight space operations directly manipulate the model parameters for various goals e.g., multi-task learning (Yu et al., 2024) and continual learning (Marouf et al., 2025; Marczak et al., 2025). For instance, weight averaging is utilized to merge the knowledge of multiple models (Ilharco et al., 2022; Jang et al., 2025), while weight reset is used to enhance plasticity (Dohare et al., 2024) and improve parameter-redundancy (Yadav et al., 2024). For an overview of weight-space operations, please refer to Section A.6.
3 Problem Formulation
Notation We generally follow the setting of class-incremental learning (class-IL) (Masana et al., 2022). We consider a sequence of $T$ tasks, each associated with distribution $P_{t}$. Let $\mathcal{D}_{t}=\{(x^{t}_{i},y^{t}_{i})\}_{i=1}^{|\mathcal{D}_{t}|}$ with $(x^{t}_{i},y^{t}_{i})\sim P_{t},$, be the training dataset for task $t$. The tasks are presented in order $t=1,\cdots,T$. The model $F$ (its parameters $\theta$) does not retain explicit access to previous task datasets $\mathcal{D}_{j}$ for $j<t$, except via a exemplar memory buffer $\mathcal{M}$ of a capacity of $K$. Thus, at the training step of task $t$, the model updates its parameters $\theta$ using the data $D_{t}\cup\mathcal{M}_{1:t-1}$ and the cross-entropy loss $\ell(\theta;x,y)$, where $\mathcal{M}_{1:t-1}$ includes selected samples from earlier tasks.
(a) Acc. graph on CIFAR-100
(b) Training time spent on CIFAR-100
Figure 1: A comparison of (a) average class-IL accuracy and (b) training time across different exemplar memory sizes in the 10 task scenario on CIFAR-100. As memory size increases, catastrophic forgetting is reduced, while training time (i.e., computational cost) increases proportionally. Note that both DER and FOSTER are expansion-based methods, represented by dashed lines, where FOSTER doubles the model size, and DER scales based on the number of tasks.
What changes under a sufficient exemplar memory regime?   We claim that a sufficiently large memory can closely approximate the data distributions of previous tasks, effectively reducing forgetting. For the complete derivation of this claim, please refer to Section A.1. Experimentally, Figure 1(a) demonstrates that a significant drop in forgetting occurs when exemplar memory is abundant. Notably, we observe that the simplest baseline (i.e., Reply) outperforms its competitors with significantly lower cost (see Figure 1(b)). This underscores two key points: First, when the sufficient exemplar memory is available, existing algorithms fall behind even the simplest baseline, making their additional computational cost unnecessary. With these observations in mind, we aim to develop a new algorithm that is (1) cost-effective and (2) addresses the new challenges that arise in a sufficient exemplar memory setting.
4 Weight Space Consolidation
1 Input: Model parameters $\theta$, training data $D_{1:t}$, memory buffer $\mathcal{M}$, average interval $j$, warming epoch $n_{\mathrm{warm}}$
2Output: Trained model parameters $\theta$
3for $t\leftarrow 1$ to $T$ do
       $\Theta\leftarrow\theta$ // Init. Averaged Model
4       if $t>1$ then
5             $\theta_{\mathrm{prev}}\leftarrow\theta$
6
7      for $i=1:n_{\mathrm{iter}}$ do
8             Sample minibatch from $\{D_{t}\cup\ \mathcal{M}_{1:t-1}\}$
9            Update $\theta$ using $\ell(\theta;x,y)$ and SGD
10            if ($t>1$ and $i=n_{\mathrm{warm}}$) then
11                   $\mathcal{I}_{\mathrm{reset}}\leftarrow\texttt{FindDormantParams}(\theta,\theta_{\mathrm{prev}})$
12                  for $l\in\mathcal{I}_{\mathrm{reset}}$ do
13                         Reset weights
14                         using Equation 1
15
16
17            if ($t>1$ and $i>n_{\mathrm{warm}}$ and i%j=0) then
18                   $n_{\mathrm{avg}}\leftarrow i/j$
19                  $\Theta\leftarrow(\Theta\cdot n_{\mathrm{\mathrm{avg}}}+\theta)/(n_{\mathrm{\mathrm{avg}}}+1)$
20
21      $\theta\leftarrow\Theta$
Algorithm 1 Weight-space consolidation for cost-efficient CL
We introduce a simple but cost-effective method that leverages weight-space operations (e.g., weight-reset, weight-average) for continual learning with sufficient exemplar memory. The grounding idea of our approach is that the sufficient exemplar memory regime allows the model to obtain an effective minimizer of all previous tasks (discussed in Section 3). Deriving from this, at the beginning of training each $t^{th}$ task, we assume that we already have a good minimizer of the previous tasks, and encourage the model to stay in its proximity in the weight-space.
Algorithm 1 shows pseudo-code of our method. Prior to the training of the $t^{th}$ task, the model parameters (i.e., minimizer of previous $11:t-$ 1 tasks) are stored as $\theta_{\textrm{prev}}$ (see Line 6 in Algorithm 1). Then, we begin training on the current $t^{th}$ task using the loss $\ell(\theta;x,y)$ (see Line 8). After $n_{\textrm{warm}}$ warming epochs, we apply a rank-based weight resetting procedure to reset the parameters that were dormant during training (see Line 11). We rank the dormant parameters by comparing the current value in $\theta$ to the value of the previous task model $\theta_{\textrm{prev}}$. We tested multiple metrics to quantify the dormancy and found that a simple distance $\big{|}\theta[l]-\theta_{\textrm{prev}}[l]\big{|}$ is sufficient to measure the drift for the $l^{th}$ parameter. Please refer to section A.3. Specifically, we reset 80% of the parameters, following empirical results of Yadav et al. (2024) (see Lines 12-14). Formally, the $l^{th}$ parameter of $\theta$ is reset using:
$\theta[l]=\alpha\cdot\theta_{\mathrm{prev}}[l]+(1-\alpha)\cdot\theta[l],$ (1)
where $\alpha=0.5$ by default. This step is intended to reset parameters that did not contribute to learning the current task $t$, allowing the model $\theta$ to remain in proximity to the previous task model $\theta_{\textrm{prev}}$. Assuming the reset model is a good (if not optimal) minimizer of the current task $t$, and also retaining the parameters of $\theta_{\textrm{prev}}$, we apply a periodic weight-averaging over the learning trajectory (see Lines 16-17) (Izmailov et al., 2018). The weight-averaging procedure is intended to accelerate convergence and improve generalization by aggregating diverse learning signals. We find this particularly beneficial under sufficient exemplar memory settings, where the variance of data is significant. Please refer to Section A.2 for a detailed elaboration of our method.
5 Experiment
Experimental settings   We evaluate our algorithm on two standard class-incremental learning (class-IL) benchmarks (Masana et al., 2022) using the PyCIL framework (Zhou et al., 2023). CIFAR100 is an image classification dataset with 100 classes, which is split into 10 sequential tasks, each containing 10 classes. ImageNet100 is another classification dataset with 100 classes, which is also split into 10 sequential tasks, each containing 10 classes. We compare our approach with six exemplar-based baselines, including iCaRL (Rebuffi et al., 2001), BiC (Wu et al., 2019), WA (Zhao et al., 2020), DER (Yan et al., 2021), and FOSTER (Wang et al., 2022). Additionally, we include Replay, a naive baseline that finetunes the model using both the current dataset and stored exemplars. All experiments use ResNet-32 for CIFAR-100 and ResNet-18 for ImageNet-100 as the backbone  (He et al., 2016). We conduct experiments across five random seeds and report the final class-IL accuracy, computed as the average classification accuracy across all tasks on the validation datasets after training on the final task (Masana et al., 2022). Experimental details are reported in Section A.5
Table 1: Average class-IL accuracy (Acc.) with different exemplar memory sizes. We report experimental results with varying memory sizes, ranging from 20 exemplars per class (a common setting in class-IL) to 500 exemplars per class (storing the entire previous dataset) from previous tasks. The bold indicates the best performance among the non-expansion-based methods.
(a) Acc. on CIFAR-100
Memory Size (the number of exemplars per class) Method 20 40 80 100 200 300 400 500 DER 64.30 67.56 70.66 71.68 74.02 74.80 75.44 75.55 FOSTER 65.83 64.95 67.00 68.67 73.46 76.01 77.39 78.24 Replay 48.92 56.59 63.30 65.12 70.84 73.46 75.38 76.90 iCaRL 49.56 57.74 64.14 65.91 71.42 73.91 75.28 76.78 BiC 54.28 60.48 63.63 65.84 68.97 70.85 72.16 72.68 WA 60.87 64.15 66.47 67.82 70.40 71.27 72.61 73.52 Ours 51.90 59.66 66.24 68.41 73.48 75.83 77.44 78.25
(b) Acc. on ImageNet-100
Memory Size (the number of exemplars per class) Method 20 50 100 200 300 400 500 600 DER 71.73 75.39 77.18 78.94 79.86 79.97 80.16 80.29 FOSTER 70.14 67.69 70.74 76.01 79.03 80.94 81.87 82.79 Replay 50.52 61.64 68.49 73.79 76.93 78.59 80.25 81.08 iCaRL 50.32 59.76 67.23 73.57 76.84 78.45 79.63 80.87 BiC 59.31 65.98 70.63 74.14 76.22 77.51 78.76 79.29 WA 61.44 67.52 71.33 75.85 77.53 78.79 79.63 80.21 Ours 54.55 65.30 71.66 76.34 78.75 80.33 81.50 82.40
(a) Training time spent on ImageNet-100
(b) Ablation Study on CIFAR-100
Figure 2: Training cost analysis and ablation study. We report experimental results with different exemplar memory sizes, ranging from 20 exemplars per class (a common setting in class-IL) to 500 exemplars per class (storing the entire previous dataset) from previous tasks.
Experimental results   Figure 2 (and Figure 1(a)) presents experimental results on the CIFAR-100 and ImageNet-100 datasets in a 10 task scenario, evaluating various exemplar memory sizes. To ensure a fair comparison, we distinguish expansion-based methods, such as DER and FOSTER, from other algorithms, as these methods increase the model size during class-IL. Based on these results, we make the following key observations: First, in the constrained exemplar memory setting typically considered (e.g., 20 exemplars per class), existing algorithms outperform the naive Replay approach. However, as the number of exemplars per class increases, the performance gap between these algorithms and Replay narrows significantly. Notably, when the exemplar count per class reaches 100 (i.e., using sufficient exemplar memory), Replay achieves competitive performance compared to other methods. Second, Figure 1(b) and Figure 2(a), which compare training costs (measured by training time), show that as exemplar memory size increases, the training cost also rises. In particular, expansion-based methods achieve superior performance but require 2-3 times the training cost of Replay. In contrast, note that our weight space consolidation method achieves superior performance across both benchmarks, nearly matching the training cost of the simplest Replay method. We also report the results of an ablation study, which demonstrates that each component of our method contributes uniquely and effectively. When combined, these components yield the strongest performance (see Figures 2(b)). We provide more detailed analyses of (1) the ablation study and (2) the effect of weight space operations on continual learning in Section A.4.
6 Concluding Remarks
Our work shifts the focus in continual learning from memory efficiency to computational cost, demonstrating that with sufficient exemplar memory, exemplar-based methods like Replay can achieve competitive performance at a fraction of the training cost. We introduce a simple, effective method that directly manipulates the model’s weight space, outperforming existing algorithms in terms of both accuracy and efficiency. These results highlight the potential for optimizing computational efficiency in class-incremental learning and provide a promising direction for practical applications, suggesting that exemplar memory constraints may no longer be the primary bottleneck in many real-world scenarios.
Appendix A Appendix
A.1 An investigation into the sufficient exemplar memory regime
Defining the sufficient exemplar memory regime
Most prior work assumes a strict exemplar memory constraint: $K\ll\sum_{t=1}^{T}|\mathcal{D}_{t}|$. Under such limited capacity, the replay buffer $\mathcal{M}$ can only hold a small fraction of each older dataset in $\{\mathcal{D}_{1},\cdots,\mathcal{D}_{t-1}\}$. Consequently, it provides only a partial representation of the task distributions $\{P_{1},\cdots,P_{t-1}\}$, leaving the model prone to catastrophic forgetting (Kirkpatrick et al., 2017) between tasks (i.e., inter-task forgetting)
By contrast, we say the replay buffer $\mathcal{M}$ is sufficient if it can store enough samples to approximate each past distribution $P_{j}$ for $1\leq j<t$. Formally, we let $K\approx\kappa\sum_{j=1}^{t-1}|\mathcal{D}_{j}|,$ where $0<\kappa\leq 1$ determines what fraction of the earlier task data can be retained. A larger $\kappa$ indicates that $\mathcal{M}$ preserves a representative set of samples from past tasks - though not necessarily every sample. As a result of this sufficient exemplar memory setting, the mixture distribution $P_{past}$ of previously encountered tasks at the training step of task $t$ effectively becomes:
$P_{past}(x,y)\approx\frac{\sum_{j=1}^{t-1}\pi_{j}P_{j}(x,y)}{\sum_{j=1}^{t-1}\pi_{j}},$ (2)
where $\pi_{j}$ denotes the relative importance (i.e., size) of each past task and $P_{j}(x,y)$ the true data distribution of task $j$. Hence, when learning task $t$, the effective training distribution is
$P_{train}^{(t)}(x,y)\approx\alpha_{t}P_{t}(x,y)+(1-\alpha_{t})P_{past}(x,y),$ (3)
where $\alpha_{t}\in[0,1]$ is a balancing factor that weights the newly introduced and replayed data, usually emphasizing the current task.
Sufficient exemplar memory mitigates catastrophic forgetting
(a) CIFAR-100
(b) ImageNet-100
Figure 3: The impact of exemplar memory size on catastrophic forgetting. Increased memory drastically reduces the forgetting between tasks, while it persists.
Next, we discuss how a sufficiently large exemplar memory buffer $\mathcal{M}$ approximates previous tasks’ distributions, thus reducing inter-task forgetting.
Ideally, if we could train jointly on all tasks $1,\cdots,t$, the obtained model parameter $\theta^{*}_{1:t}$ would minimize the risk:
$R_{1:t}(\theta)=\mathbb{\sum}_{j=1}^{t}\mathbb{E}_{(x,y)\sim P_{j}}\bigr{[}\ell(\theta;x,y)\bigr{]},$ (4)
which minimizes inter-task forgetting by design. By joint-training on all task data up to task $t$, we would find $\theta^{*}_{1:t}=\operatorname{arg}\operatorname{min_{\theta}R_{1:t}(\theta)},$ with no conventional notion of forgetting between tasks.
Similarly, with a large enough exemplar memory, the replayed data for previous tasks closely approximates their true distributions ($\tilde{P}_{j}\approx P_{j}$ for all $j<t$). Therefore training on $D_{t}\cup\mathcal{M}_{1:t-1}$ yields a risk
$\tilde{R}_{1:t}(\theta)\approx\alpha_{t}\mathbb{E}_{(x,y)\sim P_{t}}\bigr{[}\ell(\theta)\bigr{]}+(1-\alpha_{t})\mathbb{\sum}_{j=1}^{t-1}\hat{\pi}_{j}\mathbb{E}_{(x,y)\sim\tilde{P_{j}}}\bigr{[}\ell(\theta)\bigr{]},$ (5)
where we may bound $|\tilde{R}_{1:t}(\theta)-R_{1:t}(\theta)|\leq\epsilon$ under standard assumptions (e.g., Lipschitz continuity in $\theta$ (Khromov & Singh, 2024)), with $\epsilon$ shrinking as the replay buffer size $K$ increases, leading to $\tilde{P_{j}}\approx P_{j}$. A small gap between $\tilde{R}_{1:t}(\theta)$ and $R_{1:t}(\theta)$ implies that $\lVert\tilde{\theta}^{*}_{1:t}-\theta^{*}_{1:t}\rVert$ is also small, where $\tilde{\theta}^{*}_{1:t}=\operatorname{arg}\operatorname{min}_{\theta}\tilde{R}_{1:t}$. Intuitively, if the two risk surfaces are proximate, their minimizes are also close in the parameter space (Fornasier et al., 2021).
In this sense, inter-task forgetting on a previous task $j$ arises when $\tilde{\theta}^{*}_{1:t}$ drastically changes predictions on the previous task distribution $P_{j}$. But if $\tilde{\theta}^{*}_{1:t}$ remains near $\theta^{*}_{1:t}$ (which, by definition, does well on previous tasks by training jointly), it must still perform well on task $j$. Hence, if we measure inter-task forgetting in the parameter space as
$\Delta_{j\rightarrow t}=\mathbb{E}_{x\sim P_{j}}\bigr{[}\ell(\tilde{\theta}^{*}_{1:t};x)-\ell(\theta^{*}_{1:j};x)\bigr{]},$ (6)
the measure of how replay-based parameters after task $t$ perform on task $j$ compared to the parameters after task $j$. Here, $\Delta_{j\rightarrow t}$ remains small if $\tilde{\theta}^{*}_{1:t}\approx\theta_{1:t}^{*}$. Naturally, since $\theta^{*}_{1:t}$ is a reliable minimizer on all tasks, the small parameter drift ensures that the performance of the model trained under sufficient exemplar memory does not degrade on earlier tasks - i.e., inter-task forgetting is reduced as exemplar memory becomes larger (Merlin et al., 2022; Brignac et al., 2023).
New challenges under sufficient exemplar memory
While the sufficient exemplar memory regime mitigates inter-task forgetting by closely approximating past data distributions, it also creates a hybrid training dataset that combines current task data $D_{t}\cup\mathcal{M}_{1:t-1}$ with replayed samples from previous tasks.
This combined dataset introduces several challenges. First, gradient interference may occur as the gradients computed on current task samples can conflict with those from past tasks, leading to partial cancellation of updates and impeding effective learning across tasks. We view that this potentially could lead to a new form of forgetting that distorts the learned features and ultimately interfering with the model’s learning process. Second, the heterogeneous nature of the hybrid dataset increases the variance of the stochastic gradient estimates, resulting in slower convergence. This heightened variance necessitates either more iterations or more sophisticated optimization techniques to minimize the loss reliably. Furthermore, imbalances in task representation can arise if the replay buffer unevenly captures the diverse distributions of past tasks.
Conventional continual learning methods typically address forgetting by assuming that new data is markedly different from prior data and focusing on preserving performance on previously learned tasks. However, in the sufficient exemplar memory regime, where all tasks are presented simultaneously within a combined dataset, these methods fall short. They are not equipped to handle the multi-task learning dynamics and the associated balancing issues that emerge when the model must integrate and harmonize learning signals from a diverse set of tasks. We believe a more thorough investigation is required, and we set this as a key objective of our future work.
A.2 Method (continued.)
In this section, we elaborate on our method, highlighting the role of each component. The proposed method is a combination of two weight-space operations. (1) Weight Reset (Yadav et al., 2024) and (2) Weight Averaging (Wortsman et al., 2022).
Figure 4: Weight Reset Experiment on CIFAR-100. Most parameters are redundant in learning new tasks.
Weight Reset
The weight reset step selects and resets parameters that were redundant in adapting to the new task. The aim of this procedure is to find the minimal set of parameters that are capable of adapting to the new $t^{th}$ task, and reverting the redundant parameters to the value of the previous task model $\theta_{prev}$, which helps recover learned features. In this process, we use a simple distance metric to measure a model’s contribution to the new task, formulated as:
$\big{|}\theta[l]-\theta_{prev}[l]\big{|},$ (7)
where $\theta[l]$ and $\theta_{prev}[l]$ indicates the $l^{th}$ parameter of the current model $\theta$ and the previous task model $\theta_{prev}$, respectively. Using this metric, we retain the top-$Q\%$ parameters that have largely drifted during the current task and reset the dormant parameters using eq. 1. The idea of resetting model weights is not novel in the continual learning literature, but most focused on improving plasticity (Dohare et al., 2024). Conversely, we improve both plasticity and stability using a mixing technique (eq. 1, integrating previous and present tasks. In section A.4, we provide experimental results on the effect of $Q$, the percentage of weight reset on model performance (see fig. 4).
Weight Average
The weight-averaging step aggregates various learned signals into a single model, allowing for faster convergence and improved generalization (Izmailov et al., 2018). The underlying idea is that the model weights can address catastrophic forgetting, functioning as a substitute for replay memory. Using this, we aim to fill the gap between the full exemplar memory setting (i.e., joint-training) and the sufficient exemplar memory setting. Another motive behind this approach is the new challenges (see section A.1) rooted in training with a mixture of datasets $D_{t}\cup\mathcal{M}_{1:t-1}$, which generally requires more training steps for convergence. In Section A.4, we empirically show that the sufficient exemplar memory setting complicates training, requiring more training epochs for convergence (see Figure 5). The weight averaging technique is also an emerging practice in the continual learning literature (Marouf et al., 2025; Kozal et al., 2024). However, such works merge the model weights after training, which requires the storage of multiple model weights (proportional to the number of tasks). On the other hand, our method uses a moving averaged model that is updated during training, reducing the cost of storage.
A.3 Measuring Parameter-Drift
In our work, we measure a parameter’s contribution to learning a new task by comparing its parameter drift (eq. 7). However, there are a number of alternative approaches we could take. For instance, we can differentiate between how a parameter changes (1) between tasks (2) within a task.
Inter-task Parameter Drift
When transitioning from one task to the next, the change in the $l$-th parameter can be measured as
$\Delta\theta_{\text{inter}}[l]=\bigl{|}\theta_{t}[l]-\theta_{t-1}[l]\bigr{|},$
where $\theta_{t}[l]$ denotes the $l^{th}$ parameter after training on the current task $t$, and $\theta_{t-1}[l]$ represents the $l^{th}$ parameter after training on the previous task $t-1$. A large value indicates that the parameter is highly task-specific, while a small value suggests robustness across tasks.
Intra-task Parameter Drift
Alternatively, we can analyze how a parameter evolves during the training process of a single task. Let $\theta^{(i)}[l]$ denote the value of the $l^{th}$ parameter at the $i^{th}$ iteration during training. Then, the intra-task parameter drift is given by
$\Delta\theta^{(i)}_{\text{intra}}[l]=\bigl{|}\theta^{(i+1)}_{t}[l]-\theta^{(i)}_{t}[l]\bigr{|}.$
This measure captures the incremental updates of the parameter as the model optimizes its performance on the current task. By comparing both the inter-task and intra-task parameter changes, we gain a more comprehensive understanding of the role each parameter plays in adapting to new tasks as well as the dynamics of learning within a single task. In our future work, we will seek a more reliable metric to express a parameter’s behavior in the weight space.
A.4 Experimental Results and Analysis
In this section, we continue our discussion of the experimental results.
Table 2: Ablation study on CIFAR-100
Memory Size Method 20 40 80 100 200 300 400 500 Replay 48.92 56.59 63.30 65.12 70.84 73.46 75.38 76.90 w/o reset 50.23 58.19 65.01 66.50 72.33 75.00 76.98 77.92 w/o avg. 48.73 56.89 63.43 65.22 70.81 73.49 74.99 76.47 Ours 51.91 59.66 66.24 68.41 73.48 75.83 77.44 78.25
We begin with the ablation study. In this experiment, we perform an ablation study on the two components that our method utilizes, (1) the weight reset and (2) the weight average procedure. In table 2, we report the results of the experiment. Here, w/o reset indicates the replay method combined with weight average (hence without reset), while w/o avg. indicates the replay method added with the reset procedure (i.e., without avg.). The replay method is displayed to show the case when neither weight reset nor weight average is utilized. Our method showed the strongest results compared to its ablations. Notably, we can see that weight reset alone (w/o avg.) does not show strong gains. However, when the weight reset is combined with the weight average method, it allows significant performance gains.
Another experimental result is the weight reset experiment in fig. 4. Here, we can see that approximately only $20\%$ of the model parameters are responsible for the model’s learning process (i.e., resetting $80\%$ of the parameters does not pose significant issues). In contrast, when the percentage drops to $10\%$ or less, the model’s task accuracy drops by a large margin. We further discover that the weight average step helps the model to become even more sparse (i.e., not reliant on individual parameters), improving the model performance under extreme resetting settings.
Table 3: Average Acc. on CIFAR-100 with different exemplar memory sizes (full)
Memory Size Method 20 40 60 80 100 120 140 160 180 200 220 240 260 280 300 350 400 450 500 DER 64.30 67.56 69.64 70.66 71.68 73.05 73.37 73.41 73.52 74.02 74.85 75.15 74.74 74.96 74.80 75.56 75.44 75.77 75.55 FOSTER 65.83 64.95 65.36 67.00 68.67 69.85 70.84 72.48 73.33 73.46 73.87 74.63 75.16 75.96 76.01 76.88 77.39 77.67 78.24 Replay 48.92 56.59 60.78 63.30 65.12 66.67 68.00 69.23 70.10 70.84 71.61 72.16 73.10 73.19 73.46 74.643 75.38 75.85 76.90 iCaRL 49.56 57.74 61.57 64.14 65.91 67.54 69.21 69.86 70.43 71.42 71.71 72.31 72.83 73.57 73.91 74.85 75.28 76.19 76.78 BiC 54.28 60.48 62.47 63.63 65.84 67.18 67.87 68.19 68.77 68.97 69.46 69.72 70.15 70.34 70.85 71.67 72.16 72.95 72.68 WA 60.87 64.15 65.36 66.47 67.82 68.42 69.09 69.85 70.13 70.40 70.57 71.04 71.11 71.24 71.27 71.92 72.61 73.35 73.52 Ours (w/o avg.) 50.23 58.19 62.12 65.01 66.50 68.25 69.69 70.85 71.66 72.33 72.64 73.78 74.1 74.81 75.00 75.94 76.98 77.37 77.92 Ours (w/o reset) 48.73 56.89 60.81 63.43 65.22 66.50 68.28 68.84 69.85 70.81 71.29 72.03 72.77 73.50 73.49 74.68 74.99 75.57 76.47 Ours 51.91 59.66 63.88 66.24 68.41 69.96 70.92 71.99 72.67 73.48 74.11 74.80 74.87 75.65 75.83 76.89 77.44 77.95 78.25
Figure 5: The training loss under a sufficient exemplar memory continual learning setting. When training with sufficient exemplar memory, convergence becomes difficult.
Lastly, we show that under sufficient exemplar memory, convergence becomes difficult, as discussed in section A.1. The results are illustrated in Figure 5. Specifically, we analyze the training loss of the model under two cases. (1) Continual Learning: train the model across 20 tasks sequentially, with sufficient exemplar memory. (2) Joint Training: train individual models for each task using all known task data. For this experiment, we used the CIFAR-100 dataset divided into 20 subtasks. In Figure 5, we can see that under sufficient exemplar memory, converging to the training data becomes difficult, especially for the continual learning model. While a simple solution would be to extend the training epochs for convergence, it would collide with our aim for cost-efficiency. On the other hand, joint-trained models relatively converge better. This result aligns with the results of Dohare et al. (2024) that a model sequentially trained on different tasks (i.e., continual learning) suffers a loss of plasticity (i.e., a model’s ability to learn new tasks). We believe a thorough investigation of this new phenomenon is required.
A.5 Experimental Details
In this section, we report the experimental details of our experiments. In all our experiments, we have used the PyCIL (Zhou et al., 2023) library, which allows easy replication. We followed the standard training hyperparameters of the PyCIL library, which are fixed across experiments. Regarding unique hyperparameters, the average interval $j$ was set as $5$, and the warming epoch $n_{warm}$ was set as $25\%$ of the total training epochs (default set as $70$ under the PyCIL setting). $j$ and $n_{warm}$ were selected using a grid search. Please refer to Section 4 for a better understanding of each hyperparameter. Lastly, regarding the computing resources, we used a single NVIDIA RTX 6000 GPU for all experiments. For our experiments, we used the 2.2.1 version of Pytorch (Paszke et al., 2019).
A.6 Related Works (continued.)
Weight Space Operations
Recent works have shown that manipulating model parameters directly with weight-space operations (e.g., model merging (Wortsman et al., 2022)) can handle multi-task learning Yu et al. (2024) and continual learning (Marouf et al., 2025; Marczak et al., 2025) in a more principled way. These techniques usually intervene post-training by merging the weights of different models e.g., Yadav et al. (2024) suggested a selective merging algorithm that mitigates the interference between different models, while Ilharco et al. (2022) showed that arithmetic operations on the weight space can edit the model without training. Unlike these post-training interventions, our approach manipulates the model’s weight space amidst training (Izmailov et al., 2018; Jang et al., 2025) without storing multiple model parameters, aiming for cost-effective editing of the continual learner. Another relevant yet overlooked topic is the effect of weight-space operations on model attributes e.g., loss landscape (Li et al., 2018; Kaur et al., 2023) and plasticity (Dohare et al., 2024), that contribute to continual learning and generalization. This work empirically investigates various aspects of the model to study their effect on the model’s ability to handle distribution shifts. In the continual learning literature, several works have adopted weight-space operations to obtain multi-task solutions without joint training. For instance, Kozal et al. (2024) has suggested the use of weight averaging techniques for continual learning, and Marczak et al. (2025) has extended the idea using task arithmetic. However, these approaches merge models post-training and requires the storage of multiple model weights during training. On the other hand, our approach utilizes weight-space operations amidst training, without the redundant storage of multiple model weights. We view this as an important difference in modern settings where the model size is exponentially growing.